# Tight Bounds for Sparsifying Random CSPs

### v1

@@ Abstract
In order to understand the “typical” bottlenecks for CSP sparsification, we initiate the study of the sparsification of random CSPs. In particular, we consider two natural random models: the $r$-partite model and the uniform model. In the $r$-partite model, CSPs are formed by partitioning the variables into $r$ parts, with constraints selected by randomly picking one vertex out of each part. In the uniform model, $r$ distinct vertices are chosen at random from the pool of variables to form each constraint.
@@
In the uniform model, the situation is a bit more complex. For every predicate $P$, there is an integer $k$ such that a random instance on $n$ vertices and $m$ edges cannot (essentially) be sparsified if $m\leq n^{k}$ and can sparsified to size $\approx n^{k}$ if $m\geq n^{k+1}$. However, for some predicates $P$, if $m\in[n^{k},n^{k+1}]$, there may or may not be a nontrivial sparsifier. In fact, we show that there are predicates where the sparsifiability of random instances is non-monotone, i.e., as we add more random constraints, the instances become more sparsifiable. We give a precise (efficiently computable) procedure for determining which situation a specific predicate $P$ falls into.
@@ 1.1 Background
Sparsification is a powerful algorithmic technique used to reduce the size of a combinatorial object while still providing strong theoretical guarantees on the similarity of the resulting sparsified object to the original object. Since its introduction in the seminal work of Benczúr and Karger [BK96] in the context of cut-sparsification of graphs, sparsification has developed into a rich, multi-faceted field, with works studying various generalizations of cut sparsification like spectral sparsification [BSS09, ST11], hypergraph sparsification [KKTY21b, KKTY21a, Lee23, JLS23], and code sparsification [KPS24a, KPS25a, BG25], as well as algorithms for designing sparsifiers in restricted, sublinear models of computation like streaming and fully dynamic algorithms [AGM12a, AGM12b, KLM+14, ADK+16, KPS24b].
Recently, a prominent focus in the study of sparsification has shifted to the framework of CSP sparsification, as first introduced in the work of Kogan and Krauthgamer [KK15]. In this setting, one is provided with a constraint satisfaction problem (CSP), specified by a universe of variables $x_{1},\dots x_{n}$, $x_{i}\in D$, a relation $R\subseteq D^{r}$ (sometimes referred to as a predicate), and the constraints themselves, which are the result of applying the relation $R$ to ordered subsets of $r$ variables, along with a non-negative weight. Formally, the $i$th constraint is the relation $R$ applied to $S_{i}\in P([n],r)$ (namely, all ordered subsets of size $r$) with weight $w_{i}\in\mathbb{R}^{\geq 0}$, and is denoted by $C_{i}=w_{i}\cdot R_{S_{i}}(x)$, where we understand $R_{S_{i}}(x)=1$ if and only if $x_{S_{i}}\in R$, and is otherwise 0.111In general, a well-known reduction allows us to also consider CSPs with constraints that use different relations. For such a CSP, we can simply decompose it into independent instances, where each new instance contains all constraints using a single relation. We can then sparsify these instances separately and take their union to get a sparsifier of the original CSP instance. In general then, for an assignment $x\in D^{n}$ to the $n$ variables, we say that the value taken by the CSP on assignment $x$ is exactly
@@
Due to the flexibility in choosing the relation $R$, CSP sparsification captures a wide variety of natural sparsification tasks. For instance, when $R=\{01,10\}\subseteq\{0,1\}^{2}$, CSP sparsification is equivalent to cut-sparsification in graphs, when $R=\{0,1\}^{r}-\{0^{r},1^{r}\}$, CSP sparsification captures hypergraph sparsification for hyperedges of size $r$, and when $R=\{x\in\{0,1\}^{r}:\mathrm{Ham}(x)\mod 2=1\}$, CSP sparsification captures so-called code sparsification over $\mathbb{F}_{2}$. [^2] Specifically, this will capture codes whose generating matrices have $\leq r$ 11’s per row. As the relation $R$ varies more generally, CSP sparsification instead captures a large sub-class of generalized hypergraph sparsification problems as introduced in Veldt, Benson, and Kleinberg [VBK22] and which have found numerous applications. [^3] See [VBK22] for the many applications of this technique.
@@
However, this characterization in terms of the non-redundancy is still lacking in the following important respects:
@@
    In general, non-redundancy of arbitrary relations is not known to be efficiently computable, and in fact the asymptotic behavior of the non-redundancy of even some concrete small relations remains a mystery (c.f. [BCK20, BG25, BGJ+25]). Thus, neither computing the optimal worst-case sparsifier size nor the actual computation of the sparsifier itself are known to be doable in polynomial time..
@@
    Non-redundancy is only known to provide a worst-case guarantee on the sparsifiability of CSPs with a given relation. For any fixed CSP instance, the non-redundancy does not necessarily govern the sparsifiability.
In practice, many of the CSPs we encounter are not worst-case, and thus one may hope that a more efficient characterization is possible for typical instances of CSPs. In fact, such hopes were recently expressed by Brakensiek and Guruswami [BG25] where they explicitly posed the question of understanding this typical behavior. Towards this end, in this work we consider the following guiding question:
What governs the sparsifiability of random CSPs?
@@ 1.2 Our Results
At face value, it may be unclear that in either of these models, random CSPs should be more sparsifiable than their worst-case counterparts; indeed, as the value of $m$ approaches $n^{r}$, then virtually every constraint will be included in the random CSP instance $C$ that we generate. In particular then, any worst-case CSP instance is indeed a sub-instance of our CSP $C$, and so the actual non-redundancy of $C$ will be as large as the worst-case sparsifiability of CSPs with the relation $R$.
However, it turns out that the presence of these unsparsifiable sub-instances is not an inherent barrier towards sparsification of the global instance. For illustration, one can consider the relation $R=\{01\}\subseteq\{0,1\}^{2}$, which exactly models the directed-cut function in graphs. In the worst-case, there is a well known lower bound on the sparsifiability of directed graphs, as one can show that the directed graph $G=(L\sqcup R,E=L\times R)$ requires all $|E|=|L|\times|R|=\Omega(n^{2})$ directed edges to be kept (see [CCPS21] for a more detailed argument). But, what then is the sparsifiability of random directed graphs?
It turns out that random directed graphs almost always admit $(1\pm\varepsilon)$ cut-sparsifiers of size $\widetilde{O}(n/\varepsilon^{2})$. One way to see this is that random directed graphs are almost Eulerian (with high probability), meaning that their out-degree minus in-degree is small in absolute value for all/most vertices, and thus can essentially be modeled by undirected graphs, for which we do know near-linear size sparsifiability results (a formal version of this argument appears in [CCPS21]).
@@
    There exists a constant $\varepsilon>0$ for which $C$ does not admit a $(1\pm\varepsilon)$ sparsifier of size $\Omega(\min(m,n^{c}))$.
Thus, in the $r$-partite setting, the sparsifiability of a random CSP undergoes a strict phase transition: when the number of constraints $m$ is smaller than roughly $n^{c}$, there is no sparsification that can be done. However, as soon as the number of constraints exceeds $n^{c}$, then the CSP can be sparsified back down to $n^{c}$ constraints with high probability. In a strong sense then, this “presence of an AND” in $R$ strongly dictates the sparsifiability of random CSPs with the relation $R$. Another important property of the characterization above is its “monotonicity”: as one might expect, adding more random constraints makes sparsifying the CSP harder up to a point (roughly $n^{c}$ constraints), at which point the sparsifier size stagnates. Importantly, adding more random constraints does not decrease the sparsifier size.
@@
    If $R$ is precisely $k$-plentiful, and the symmetrization of $R$ is not marginally uniform, then (1) $C$ is sparsifiable to $\widetilde{O}_{\mathbf{R}}(\min(m,n^{r-k+1}/\varepsilon^{2}))$ constraints, and (2) for a sufficiently small, constant $\varepsilon$, any sparsifier of $C$ requires $\Omega_{\mathbf{R}}(\min(m,n^{r-k+1}))$ constraints.
@@
    If $R$ is precisely $k$-plentiful, and $R$ is marginally uniform, then (1) $C$ is sparsifiable to $\widetilde{O}_{\mathbf{R}}(\min(m,n^{r-k}/\varepsilon^{3}))$ constraints, and (2) for a sufficiently small, constant $\varepsilon$, any sparsifier of $C$ requires $\Omega_{\mathbf{R}}(\min(m,n^{r-k}))$ constraints.
@@
    If $R$ is precisely $k$-plentiful, the symmetrization of $R$ is marginally uniform, but $R$ itself is not marginally uniform, then (1) if $m\geq n^{r-k+1}$, $C$ is sparsifiable to $\widetilde{O}_{\mathbf{R}}(n^{r-k}/\varepsilon^{3})$, but (2) if $m\leq n^{r-k+1}$, there is a small, constant $\varepsilon$, such that any sparsifier of $C$ requires $\Omega_{\mathbf{R}}(m)$ constraints.
@@
In the first two cases, the CSP behaves in a similar manner to Theorem 1.3, as there is a single threshold at which the sparsifiability begins to stay the same as one adds in more constraints. However, of key importance is the distinction in the final case above. Here, for any $m\leq n^{r-k+1}$ constraints, $C$ is essentially unsparsifiable, as any sparsifier requires $\Omega_{\mathbf{R}}(m)$ constraints. However, once $m$ reaches roughly $n^{r-k+1}$, the CSP suddenly becomes more sparsifiable, with sparsifiers being computable of size roughly $n^{r-k}$. Note that this is not simply a figment of our terminology, and there are in fact concrete relations $R\subseteq D^{r}$ of constant size for which this behavior appears.
Several remarks are in order. First, it is worth noting that the above characterizations of sparsifiability depend only on the relation $R$, and so are computable in constant time (as we deal only with relations $R$ of arbitrary constant size). This, in combination with the fact that our cases are exhaustive, yields an efficient, complete characterization of the sparsifiability of CSPs in both of our studied models.
Second, and perhaps surprisingly, in both the random uniform and random $r$-partite cases, the random sparsifiability of a CSP on $n$ variables always behaves as $n^{c}$ where $c$ (which depends on the case and the relation) is integral. [^7] Of course, ignoring cases where no sparsification is possible, and $m$ itself happens to be $n^{\gamma}$, for $\gamma$ non-integral. This stands in stark contrast to the sparsifiability of worst-case instances, where it had been conjectured that the sparsifiability exponent was always integral [FK17, BŽ20, KPS24a, KPS25a], until the work of [BG25], who constructed explicit CSP instances whose sparsifiability behaves as $n^{\gamma}$ for $\gamma\notin\mathbb{Z}$. Very recently, it was even shown that every rational $\gamma\in\mathbb{Q}\cap[1,\infty)$ can appear as the exponent for some concrete CSP predicate [BGJ+25].
@@
Thirdly, in many senses our results present a very strict phase transition in the sparsifiability of random CSPs: for instance, in the $r$-partite model, our results indicate that there is a constant $c$ which depends on the relation at hand such that (with high probability) any random CSP on $n$ variables with $\leq n^{c}$ constraints cannot be sparsified, while any CSP with more than $n^{c}$ constraints can be sparsified, with the number of constraints being reduced all the way down to roughly $n^{c}$. In this manner, our results can be seen as an extension of a long line of works studying phase transitions in random CSPs (see for instance [FK96]). Perhaps more exciting though, our results also indicate a slight deviation from the typical setting of phase transitions. Indeed, random CSPs often portray monotone properties, where as one increases the number of constraints, a property (such as say, satisfiability [DSS15]) becomes strictly less likely. In our setting however, we show that there are some CSPs where the property is non-monotone, as adding more constraints makes the problem of sparsification harder up until a point, where there is then a phase transition, subsequent to which the problem becomes significantly easier.
Finally, in addition to our characterizations of the sparsifiability of random CSPs with relations $R\subseteq D^{r}$, we also consider the sparsifiability of random CSPs with valued relations $R:D^{r}\rightarrow[0,1]$, marking further progress towards understanding the sparsifiability of VCSPs as first posed in [BŽ20]. For uniform random CSPs, the sparsifiability mirrors exactly that of Theorem 1.4, while for $r$-partite random CSPs, the formal characterization is included below:
Theorem 1.5.
[Informal] Let $R:D^{r}\rightarrow\mathbb{R}^{\geq 0}$ be a valued relation, and let $C$ be a random $r$-partite CSP with relation $R$ with $m$ constraints over $nr$ variables. Then, for $c$ being the arity of the largest $\mathbf{AND}$ that can be restricted to in $R$, if every such $\mathbf{AND}_{c}$ is uniform, let $\widehat{c}=c$, and otherwise let $\widehat{c}=c+1$. [^8] See Section 6 for a formal definition. For this $\widehat{c}$, with high probability:
  1. 1.
  2. 2.
    There exists a constant $\varepsilon>0$ for which $C$ does not admit a $(1\pm\varepsilon)$ sparsifier of size $)\Omega_{\mathbf{R}}(\min(m,n^{\widehat{c}}$)).
In the following section, we summarize some of the techniques and challenges that arise in deriving our characterization.

### v2

@@ Abstract
In order to understand the “typical” bottlenecks for CSP sparsification, we initiate the study of the sparsification of random CSPs. In particular, we consider two natural random models: the $r$-partite model and the uniform model. In the $r$-partite model, CSPs are formed by partitioning the variables into $r$ parts, with constraints selected by randomly picking one variable out of each part. In the uniform model, $r$ distinct variables are chosen at random from the pool of variables to form each constraint.
@@
In the uniform model, the situation is a bit more complex. For every predicate $P$, there is an integer $k$ such that a random instance on $n$ vertices and $m$ edges cannot (essentially) be sparsified if $m\leq n^{k}$ and can be sparsified to size $\approx n^{k}$ if $m\geq n^{k+1}$. However, for some predicates $P$, if $m\in[n^{k},n^{k+1}]$, there may or may not be a nontrivial sparsifier. In fact, we show that there are predicates where the sparsifiability of random instances is non-monotone, i.e., as we add more random constraints, the instances become more sparsifiable. We give a precise (efficiently computable) procedure for determining which situation a specific predicate $P$ falls into.
@@ 1.1 Background
Sparsification is a powerful algorithmic technique used to reduce the size of a combinatorial object while still providing strong theoretical guarantees on the similarity of the resulting sparsified object to the original object. Since its introduction in the seminal work of Benczúr and Karger [BK96] in the context of cut-sparsification of graphs, sparsification has developed into a rich, multi-faceted field, with works studying various generalizations of cut sparsification like spectral sparsification [BSS09, ST11], hypergraph sparsification [KKTY21b, KKTY21a, Lee23, JLS23], sparsifying sums of positive semi-definite matrices [BKLM25, HLM+25], and code sparsification [KPS24a, KPS25a, BG25], as well as algorithms for designing sparsifiers in restricted, sublinear models of computation like streaming and fully dynamic algorithms [AGM12a, AGM12b, KLM+14, ADK+16, KPS24b].
Recently, a prominent focus in the study of sparsification has shifted to the framework of CSP sparsification, as first introduced in the work of Kogan and Krauthgamer [KK15]. In this setting, one is provided with a constraint satisfaction problem (CSP), specified by a universe of variables $x_{1},\dots x_{n}$, $x_{i}\in D$, a relation $R\subseteq D^{r}$ (sometimes referred to as a predicate), and the constraints themselves, which are the result of applying the relation $R$ to ordered subsets of $r$ variables, along with a non-negative weight. Formally, the $i$th constraint is the relation $R$ applied to $S_{i}\in P([n],r)$ (namely, ordered subsets of size $r$) with weight $w_{i}\in\mathbb{R}^{\geq 0}$, and is denoted by $C_{i}=w_{i}\cdot R_{S_{i}}(x)$, where we understand $R_{S_{i}}(x)=1$ if and only if $x_{S_{i}}\in R$, and is otherwise 0.111In general, a well-known reduction allows us to also consider CSPs with constraints that use different relations. For such a CSP, we can simply decompose it into independent instances, where each new instance contains all constraints using a single relation. We can then sparsify these instances separately and take their union to get a sparsifier of the original CSP instance. In general then, for an assignment $x\in D^{n}$ to the $n$ variables, we say that the value taken by the CSP on assignment $x$ is exactly
@@
Due to the flexibility in choosing the relation $R$, CSP sparsification captures a wide variety of natural sparsification tasks. For instance, when $R=\{01,10\}\subseteq\{0,1\}^{2}$, CSP sparsification is equivalent to cut-sparsification in graphs, when $R=\{0,1\}^{r}-\{0^{r},1^{r}\}$, CSP sparsification captures hypergraph sparsification for hyperedges of size $r$, and when $R=\{x\in\{0,1\}^{r}:\mathrm{Ham}(x)\mod 2=1\}$, CSP sparsification captures so-called code sparsification over $\mathbb{F}_{2}$. [^2] Specifically, this will capture codes whose generating matrices have $\leq r$ 11’s per row. As the relation $R$ varies more generally, CSP sparsification instead captures a large sub-class of generalized hypergraph sparsification problems as introduced in Veldt, Benson, and Kleinberg [VBK22] and which have found numerous applications, for instance, in circuit design [AK95, Law73], scientific computing / matrix multiplication [BDKS16], and in many clustering algorithms [ALZM+05, MN12, LM17, LM18]. [^3] See [VBK22] for further discussion of the applications of this technique. In all of these applications, running time depends on the sparsity of the underlying CSP (hypergraph), and this naturally motivates the study of better sparsification algorithms.
@@
However, this characterization in terms of the non-redundancy is still lacking in many important respects:
@@
    In general, non-redundancy of arbitrary relations is not known to be efficiently computable, and in fact the asymptotic behavior of the non-redundancy of even some concrete small relations remains a mystery (c.f. [BCK20, BG25, BGJ+25]). Thus, neither computing the optimal worst-case sparsifier size nor the actual computation of the sparsifier itself are known to be doable in polynomial time.
@@
    Non-redundancy is only known to provide a worst-case guarantee on the sparsifiability of CSPs with a given relation. This means that for any fixed CSP instance, the non-redundancy does not necessarily govern the sparsifiability. In fact, the non-redundancy of a fixed CSP instance is always at least as large as the worst-case sparsifiability of any sub-instance (subset of constraints) of that instance.
In practice, many of the CSP instances we encounter are not worst-case, and thus one may hope that such instances are sparsifiable to sizes smaller than their worst-case sub-instances, thereby circumventing the non-redundancy type bounds from Brakensiek and Guruswami [BG25]. In fact, [BG25] explicitly poses this question of understanding this typical behavior of the sparsifiability of CSPs. Towards this end, in this work we consider the following guiding question:
What governs the sparsifiability of random CSP instances?
Our primary contribution is to give an exact answer to the above question by providing tight bounds on the sparsifiability of random CSPs. In order to bypass the “worst sub-instance sparsifiability” lower bound that confounds previous CSP sparsifications [KPS24a, KPS25a, BG25], we introduce a new framework of decomposing and sandwiching CSPs that may be interesting in its own right.
@@ 1.2 Our Results
At face value, it may be unclear that in either of these models, random CSPs should be more sparsifiable than their worst-case counterparts; indeed, as the value of $m$ approaches $n^{r}$, then virtually every constraint will be included in the random CSP instance $C$ that we generate. In particular then, any worst-case CSP instance is indeed contained in our CSP CC!
For such CSPs, existing non-redundancy type characterizations ([BG25], and also the non-redundancy-esque characterizations of [BKLM25]) will always report a sparsifiability value which is equal to the worst sparsifiability of any sub-instance. However, it turns out that the presence of these unsparsifiable sub-instances is not an inherent barrier towards sparsification of the global instance. For illustration, one can consider the relation $R=\{01\}\subseteq\{0,1\}^{2}$, which exactly models the directed-cut function in graphs. In the worst-case, there is a well known lower bound on the sparsifiability of directed graphs, as one can show that the directed graph $G=(L\sqcup R,E=L\times R)$ requires all $|E|=|L|\times|R|=\Omega(n^{2})$ directed edges to be kept (see [CCPS21] for a more detailed argument). But, what then is the sparsifiability of random directed graphs?
It turns out that random directed graphs almost always admit $(1\pm\varepsilon)$ cut-sparsifiers of size $\widetilde{O}(n/\varepsilon^{2})$. One way to see this is that random directed graphs are almost Eulerian (with high probability), meaning that their out-degree minus in-degree is small in absolute value for all/most vertices, and thus can essentially be modeled by undirected graphs, for which we do know near-linear size sparsifiability results (a formal version of this argument appears in [CCPS21]). This argument also shows that the complete directed graph on $n$ vertices is sparsifiable to $\widetilde{O}(n/\varepsilon^{2})$ many edges. On the other hand, plugging this instance into any non-redundancy based sparsification framework will only guarantee sparsifiability to size $O(n^{2})$, as the aforementioned worst sub-instance is the complete bipartite directed graph, which is indeed contained in the complete graph.
Random $r$-Partite CSPs
@@
    There exists a constant $\varepsilon>0$ for which any $(1\pm\varepsilon)$ sparsifier of $C$ requires size $\Omega(\min(m,n^{c}))$.
Thus, in the $r$-partite setting, the sparsifiability of a random CSP undergoes a strict phase transition: when the number of constraints $m$ is smaller than roughly $n^{c}$, there is no sparsification that can be done. However, as soon as the number of constraints exceeds $n^{c}$, then the CSP can be sparsified back down to $n^{c}$ constraints with high probability. In a strong sense then, this “presence of an AND” in $R$ strongly dictates the sparsifiability of random CSPs with the relation $R$.
Note that in prior works on sparsification, the notion of the largest $\mathbf{AND}$ that can be projected to in a predicate was conjectured to govern worst-case sparsifiability [BŽ20, KPS24a, KPS25a], but this conjecture was proven false in [BG25]. Even so, these notions of projection to $\mathbf{AND}$ and restriction to $\mathbf{AND}$ are distinct, as projection allows for setting variables equal to one another (and to $\{0,1\}$), while restriction only allows setting variables to $\{0,1\}$ (and in particular, the largest $\mathbf{AND}$ under projections is always at least as large as the largest $\mathbf{AND}$ under restrictions). Nevertheless, we view it as an important contribution to now establish a formal connection between sparsifiability of CSPs and the presence of $\mathbf{AND}$’s in the relation.
Another important property of the characterization above is its “monotonicity”: as one might expect, adding more random constraints makes sparsifying the CSP harder up to a point (roughly $n^{c}$ constraints), at which point the sparsifier size stagnates. Thus, adding more random constraints does not decrease the sparsifier size.
Random $r$-Partite Valued CSPs
In addition to our characterizations of the sparsifiability of random CSPs with relations $R\subseteq D^{r}$, we also consider the sparsifiability of random CSPs with valued relations $\mathbf{R}:D^{r}\rightarrow\mathbb{R}^{\geq 0}$. Although Butti and Živný [BŽ20] make a brief remark concerning the sparsification of valued CSPs (VCSPs), we are the first work to formally do so. For $r$-partite random VCSPs, the formal characterization is included below.
Theorem 1.4.
[Informal] Let $R:D^{r}\rightarrow\mathbb{R}^{\geq 0}$ be a valued relation, and let $C$ be a random $r$-partite CSP with relation $R$ with $m$ constraints over $nr$ variables. Then, for $c$ being the arity of the largest $\mathbf{AND}$ that can be restricted to in $R$, if every such $\mathbf{AND}_{c}$ is balanced, let $\widehat{c}=c$, and otherwise let $\widehat{c}=c+1$. [^7] See Section 6 for a formal definition. For this $\widehat{c}$, with high probability:
  1. 1.
  2. 2.
    There exists a constant $\varepsilon>0$ for which any $(1\pm\varepsilon)$-sparsifier of $C$ requires size $)\Omega_{\mathbf{R}}(\min(m,n^{\widehat{c}}$)).
The key distinction between Theorem 1.3 and Theorem 1.4 is that the size of the largest $\mathbf{AND}$ appearing in the valued relation no longer solely governs the sparsifiability of the resulting VCSP instance: but rather, a lack of balance. For example, it is straightforward to compute that any instance of the CSP relation $R=\{0,1\}^{r}$ can be sparsified to a single constraint (as all constraints behave identically for all assignments). However, one can show that an arbitrary function $\mathbf{R}:\{0,1\}^{r}\to\mathbb{R}^{>0}$ where all weights are nonzero but not all equal, must have at least $\Omega(n)$ constraints in every $(1\pm\varepsilon)$-sparsifier (this is what we call an imbalanced relation). Taken for a general valued relation, we have to “zoom in” on particular parts of the relation and perform this “balanced versus imbalanced test” to determine whether $\widehat{c}=c$ (looks balanced) or $\widehat{c}=c+1$ (does not).
Random Uniform CSPs
@@
To start, in the uniform setting, we require an analogous notion of the largest $\mathbf{AND}$ in a relation $R$, which we characterize via so-called plentiful-ness. However, the more interesting distinction which arises when compared to the $r$-partite setting is in the characterization of non-monotonicity.
To better explain the intuition for this non-monotonicity in the uniform setting, let us consider a relation $R\subseteq D^{r}$ and suppose that we have a huge number of constraints (approaching $\binom{n}{r}\cdot r!$). In this case, the CSP instance we are dealing with is essentially the complete instance. In the uniform model, when we have the complete CSP, this means that we have every constraint $R_{S}(x)$, for $S\in P([n],r)$. So, for every unordered subset of variables $T\in\binom{[n]}{r}$, we have a constraint for every possible permutation of the $r$ variables in $T$. Thus, sparsifying the complete instance with relation $R$ is effectively the same as sparsifying the complete instance with what we call the symmetrization of $R$ (essentially, the average over all $r!$ permutations of the variables).
This turns out to have huge ramifications for sparsification: indeed, if we revisit the examples from the $r$-partite setting, we saw that CSPs with so-called “balanced” relations like $R=\{0,1\}^{r}$ are easy to sparsify, while CSPs with “imbalanced” relations are harder to sparsify. A key subtlety in the uniform setting is that there are relations which are themselves imbalanced but which, after symmetrization, become balanced–see Remark 7.23 for examples. The upshot for random CSP sparsification is that for denser random CSPs, sparsification becomes easier, as, after symmetrization, we are dealing with balanced relations, while for less dense random CSPs, sparsification is more difficult as we are dealing with imbalanced relations. The effect is of sufficient magnitude that it leads to a non-monotone relationship between the density of the random CSP and the size of the optimal sparsifier. We present our characterization theorem as follows.
@@
    If $R$ is precisely $k$-plentiful, and the symmetrization of $R$ is not marginally balanced, then (1) $C$ is $(1\pm\varepsilon)$-sparsifiable to $\widetilde{O}_{\mathbf{R}}(\min(m,n^{r-k+1}/\varepsilon^{2}))$ constraints, and (2) for a sufficiently small, constant $\varepsilon$, any $(1\pm\varepsilon)$-sparsifier of $C$ requires $\Omega_{\mathbf{R}}(\min(m,n^{r-k+1}))$ constraints.
@@
    If $R$ is precisely $k$-plentiful, and $R$ is marginally balanced, then (1) $C$ is $(1\pm\varepsilon)$-sparsifiable to $\widetilde{O}_{\mathbf{R}}(\min(m,n^{r-k}/\varepsilon^{3}))$ constraints, and (2) for a sufficiently small, constant $\varepsilon$, any $(1\pm\varepsilon)$-sparsifier of $C$ requires $\Omega_{\mathbf{R}}(\min(m,n^{r-k}))$ constraints.
@@
    If $R$ is precisely $k$-plentiful, the symmetrization of $R$ is marginally balanced, but $R$ itself is not marginally balanced, then (1) if $m\geq n^{r-k+1}$, $C$ is $(1\pm\varepsilon)$-sparsifiable to $\widetilde{O}_{\mathbf{R}}(n^{r-k}/\varepsilon^{3})$, but (2) if $m\leq n^{r-k+1}$, there is a small, constant $\varepsilon$, such that any $(1\pm\varepsilon)$-sparsifier of $C$ requires $\Omega_{\mathbf{R}}(m)$ constraints.
@@
In the first two cases, the CSP behaves in a similar manner to Theorem 1.3, as there is a single threshold at which the sparsifiability begins to stay the same as one adds in more constraints. However, of key importance is the distinction in the final case above. Here, for any $m\leq n^{r-k+1}$ constraints, $C$ is essentially unsparsifiable, as any sparsifier requires $\Omega_{\mathbf{R}}(m)$ constraints. However, once $m$ reaches roughly $n^{r-k+1}$, the CSP suddenly becomes more sparsifiable, with sparsifiers being computable of size roughly $n^{r-k}$. Note that this is not simply a figment of our terminology, and there are in fact concrete (even non-valued) relations $R\subseteq D^{r}$ of constant size for which this behavior appears!
@@
In many senses our results present a very strict phase transition in the sparsifiability of random CSPs: for instance, in the $r$-partite model, our results indicate that there is a constant $c$ which depends on the relation at hand such that (with high probability) any random CSP on $n$ variables with $\leq n^{c}$ constraints cannot be sparsified, while any CSP with more than $n^{c}$ constraints can be sparsified, with the number of constraints being reduced all the way down to roughly $n^{c}$. In this manner, our results can be seen as an extension of a long line of works studying phase transitions in random CSPs (see for instance [FK96]). Perhaps more exciting though, our results also indicate a slight deviation from the typical setting of phase transitions. Indeed, random CSPs often portray monotone properties, where as one increases the number of constraints, a property (such as say, satisfiability [DSS15]) becomes strictly less likely. In our setting however, we show that there are some CSPs where the property is non-monotone, as adding more constraints makes the problem of sparsification harder up until a point, where there is then a phase transition, subsequent to which the problem becomes significantly easier.
For uniform random VCSPs, the sparsifiability classification mirrors exactly that of Theorem 1.5 (see Theorem 7.15). In fact, in the uniform model the abstraction of CSPs as VCSPs greatly aids in studying the classification.
Some General Remarks
In all of our main results, it is worth noting that the characterizations of sparsifiability depend only on the relation $R$, and so are computable in constant time (as we deal only with relations $R$ of arbitrary constant size). This, in combination with the fact that our cases are exhaustive, yields an efficient, complete characterization of the sparsifiability of (V)CSPs in both of our studied models.
Furthermore, and perhaps surprisingly, in both the random uniform and random $r$-partite cases, the random sparsifiability of a CSP on $n$ variables always behaves as $n^{c}$ where $c$ (which depends on the case and the relation) is integral. [^8] Of course, ignoring cases where no sparsification is possible, and $m$ itself happens to be $n^{\gamma}$, for $\gamma$ non-integral. This stands in stark contrast to the sparsifiability of worst-case instances, where it had been conjectured that the sparsifiability exponent was always integral [FK17, BŽ20, KPS24a, KPS25a], until the work of [BG25], who constructed explicit CSP instances whose sparsifiability behaves as $n^{\gamma}$ for $\gamma\notin\mathbb{Z}$. Very recently, it was even shown that every rational $\gamma\in\mathbb{Q}\cap[1,\infty)$ can appear as the exponent for some concrete CSP relation [BGJ+25].